### v1

# SVL: Spike-based Vision-language Pretraining for Efficient 3D Open-world Understanding

Abstract
Spiking Neural Networks (SNNs) provide an energy-efficient way to extract 3D spatio-temporal features. However, existing SNNs still exhibit a significant performance gap compared to Artificial Neural Networks (ANNs) due to inadequate pre-training strategies. These limitations manifest as restricted generalization ability, task specificity, and a lack of multimodal understanding, particularly in challenging tasks such as multimodal question answering and zero-shot 3D classification. To overcome these challenges, we propose a Spike-based Vision-Language (SVL) pretraining framework that empowers SNNs with open-world 3D understanding while maintaining spike-driven efficiency. SVL introduces two key components: (i) Multi-scale Triple Alignment (MTA) for label-free triplet-based contrastive learning across 3D, image, and text modalities, and (ii) Re-parameterizable Vision-Language Integration (Rep-VLI) to enable lightweight inference without relying on large text encoders. Extensive experiments show that SVL achieves a top-1 accuracy of 85.4% in zero-shot 3D classification, surpassing advanced ANN models, and consistently outperforms prior SNNs on downstream tasks, including 3D classification (+6.1%), DVS action recognition (+2.1%), 3D detection (+1.1%), and 3D segmentation (+2.1%) with remarkable efficiency. Moreover, SVL enables SNNs to perform open-world 3D question answering, sometimes outperforming ANNs. To the best of our knowledge, SVL represents the first scalable, generalizable, and hardware-friendly paradigm for 3D open-world understanding, effectively bridging the gap between SNNs and ANNs in complex open-world understanding tasks. Code is available Here.
1 Introduction
Bio-inspired Spiking Neural Networks (SNNs) offer an efficient approach to learning superior representations from sparse 3D geometric data (e.g., event streams and point clouds) [1], owing to their distinctive spike-driven nature [2] and spatio-temporal processing capabilities [3]. For instance, the Speck [4] chip uses event-by-event sparse processing to handle 3D input data, with operational power consumption as low as 0.7 mW. However, existing SNNs [5; 6; 7] exhibit a significant performance gap compared to ANNs, and remain task-dependent, lacking both generalizable representations and the ability to achieve multimodal understanding in 3D open-world scenarios.
For instance, when deploying SNNs in real-world scenarios [4], they may struggle to generalize to input data from unseen categories not present in the training set. This highlights the critical need to develop robust pretraining strategies to enhance the visual representation capabilities and adaptability of SNNs. Existing methods, such as STDP-based initialization [8] and knowledge distillation in SpikeBert and SpikeCLIP [9; 10; 11], refine spike-based representations, while SpikformerV2 and Spike-driven Transformer V3 [6; 7] employ masked image modeling to improve scalability. However, these approaches [8] either lose effectiveness as dataset complexity increases, demand substantial computational resources [7], which limits their feasibility for neuromorphic hardware deployment, or lack multimodal integration [9]. Moreover, pre-trained models often exhibit inadequate visual representation capabilities and limited transferability, restricting unified applicability to downstream tasks [6].
Another challenge is the limited availability of annotated 3D datasets, as the creation of such datasets is both labor-intensive and error-prone, rendering it often impractical for large-scale, real-world applications [12]. In response, Vision-Language Models (VLMs) [13] have been employed to explore the transfer of knowledge gleaned from extensive 2D datasets to facilitate open-world 3D understanding. However, most VLMs [14; 15; 16] depend heavily on large-scale text encoders during inference, which imposes substantial limitations on the practicality of hardware deployment.
Figure 1: Overall architecture and applications of our SVL. (a) In pretraining, we proposed Multi-scale Triple Alignment (MTA) that jointly optimizes correlation alignment across text, image, and 3D inputs. (b) For downstream tasks, we propose Re-parameterizable Vision-Language Integration (Rep-VLI) to reparameterize the text embeddings generated by the text encoder into lightweight weights, enabling efficient spike-driven inference.
In this paper, we introduce a universal Spike-based Vision-Language pretraining framework (SVL) that enhances SNNs’ capability for open-world multimodal 3D understanding while maintaining efficient spike-driven inference. As shown in Fig. 1, our framework incorporates two key innovations: (i) Multi-scale Triple Alignment (MTA), which enables label-free triplet representation learning for capturing geometric properties of 3D data, and (ii) Re-parameterizable Vision-language Integration (Rep-VLI), which facilitates efficient deployment by addressing the computational overhead of text-driven inference. Moreover, our framework uses CLIP for its strong generalization. During pre-training, we freeze CLIP and train the 3D encoder by aligning 3D features with CLIP’s textual and visual features via contrastive learning. The pre-trained 3D model can then be fine-tuned for downstream tasks. To demonstrate SVL’s efficacy, we rigorously evaluated our models on a wide range of 3D tasks, including open-world understanding and downstream applications. We also highlight SVL’s versatility through its potential in generative tasks like 3D object captioning and open-world question answering, showcasing its broad applicability in 3D vision domains. Our main contribution can be summarized as:
  • We introduce SVL, a universal spike-based vision-language pretraining framework that enables SNNs to understand open-world multimodal 3D information while maintaining efficient spike-driven inference.
  • We propose two key technical innovations: (i) Multi-scale Triple Alignment (MTA), a label-free triplet learning mechanism for capturing geometric properties of 3D data across different scales, and (ii) Re-parameterizable Vision-language Integration (Rep-VLI), which achieves lightweight inference by reducing the computational overhead of text encoder.
  • Extensive experiments on multiple benchmarks demonstrate the effectiveness of SVL, achieving state-of-the-art (SOTA) performance in both 3D open-world understanding and downstream tasks, as well as generative applications such as 3D object captioning and open-world question answering.
2 Related Works
Pretraining Algorithms of SNNs
Numerous pretraining methods are proposed for spike-based representation learning. [8] utilized spike-timing-dependent plasticity [17] to initialize SNNs, enhancing the model’s robustness and training speed. While this approach has been successful on simple datasets with shallow networks, its effectiveness diminishes as the complexity of the datasets and networks increases. To address these issues, SpikeBert and SpikeCLIP [9; 10; 11] employ a two-stage knowledge distillation process from ANNs to enhance spike-based representations for complex downstream tasks. However, these methods rely on ANN weight initialization, limiting structural flexibility. Additionally, they use LayerNorm, which hinders neuromorphic hardware deployment. SpikformerV2 and Spike-driven Transformer V3 [6; 7; 18] apply a masked image modeling approach to address the performance degradation in SNNs as the model scales up. However, they require substantial storage and computational resources, and the lack of multimodal integration, particularly language guidance, limits their effectiveness in open-world understanding tasks.
Vision-language Models (VLMs)
aim to align image and text embeddings for cross-modal transfer, with CLIP [13] being a seminal work that uses contrastive learning for zero-shot classification. Building on this foundation, subsequent methods have expanded cross-modal alignment to include other modalities. These approaches typically fall into two categories: dual-encoder and triple-encoder frameworks. Dual-encoder fine-tune both visual and textual encoders [11; 19; 20]. Triple-encoder frameworks incorporate additional modality-specific encoders [14; 15; 21], which combine triple models to achieve open-world understanding. This architecture is highly flexible, making it suitable for a variety of downstream tasks [12; 16]. However, triple-encoder frameworks still rely on large text encoders during inference, hindering hardware deployment.
Efficient 3D Recognition
is to learn better representations from sparse and irregular 3D geometric data (e.g., event and point cloud). Currently, there exist two main approaches. Point-based methods [22; 23] utilize the PointNet series to directly extract geometric features from raw 3D data and make predictions. Voxel-based methods [24] first convert the 3D data into regular voxels and then use 3D sparse convolutions for feature extraction. Nevertheless, the improved accuracy is often accompanied by increased computational costs, limiting its applicability in practical systems Hence, numerous studies [25; 26; 27] in the SNN field combine spiking neurons with point-based methods like PointNet [22] for low energy consumption edge computing. These methods are too simplistic to apply to various tasks and datasets. It was not until E-3DSNN [5] used spike sparse convolution to achieve high performance in numerous 3D tasks with spike-driven nature. Our approach is orthogonal to the aforementioned 3D encoders. Pre-training with SVL can enhance the visual extraction capabilities of the 3D encoders, thereby improving their performance in 3D recognition tasks. SVL can also endow them with the ability to achieve open-world understanding while retaining their spike-driven nature.
3 Preliminaries
Spiking Neurons
are inspired by the dynamics of biological neurons [3; 28], which are the fundamental units of Spiking Neural Networks (SNNs). Among these, the Leaky Integrate-and-Fire (LIF) neuron is the most widely used due to its balance between biological plausibility and computational efficiency [3]. We begin by translating the LIF spiking neuron into an iterative expression using the Euler method [29], which is described as follows:
$u_{i}^{(\ell)}[t+1]=h_{i}^{(\ell)}[t]+f(w^{(\ell)},x_{i}^{(\ell-1)}[t]),$ (1)
$s_{i}^{(\ell)}[t]=\Theta(u_{i}^{(\ell)}[t+1]-\vartheta),$ (2)
$h_{i}^{(\ell)}[t+1]=\beta u_{i}^{(\ell)}[t+1](1-s_{i}^{(\ell)}[t]),$ (3)
Here, $\beta$ is the time constant $t$ and $i$ represents the time step and the neuron index in the $\ell$-th layer, respectively. The weight matrix $w$ defines the synaptic connections between adjacent layers, while $f(\cdot)$ is a function that denotes operations such as convolution (Conv) or fully connected (FC). The input is represented by $x$, and ${\Theta(\cdot)}$ denotes the Heaviside step function. When the membrane potential $u$ exceeds the firing threshold $\vartheta$, the LIF neuron generates a spike, $s$. Additionally, $h$ represents the membrane potential after the spike event, which is scaled by a constant factor $\beta$.
Directly training the above LIF-based SNNs requires the use of backpropagation through time (BPTT) [29], resulting in a time complexity of $\mathcal{O}(LT)$, where $L$ and $T$ are the number of layers and time steps. This significantly increases both the training time and memory requirements. To mitigate this issue, we use the Integer LIF Spiking Neuron.
Integer LIF Spiking Neuron
is incorporated into our SVL to reduce the quantization error, training time, and memory [6; 30; 31], which allows us to rewrite Eq. (2) as:
${s}^{(\ell)}_{i}[t]=\lfloor\operatorname{clip}\{{u}^{(\ell)}[t],0,{D}^{t}\}\rceil,$ (4)
[…]
CLIP
employs a contrastive learning framework to forge associations between images and textual descriptions [13]. It leverages a vast dataset comprising 400 million image-text pairs, training an image encoder based on either a ResNet [32] or Vision Transformer [33] architecture, and a text encoder utilizing the Transformer model [34]. These encoders work together to project both images and text into a unified embedding space. This training paradigm enables CLIP to perform zero-shot classification, allowing it to identify images solely based on textual descriptions, even without category-specific training. In this study, we aim to utilize CLIP’s semantically rich feature space to support spike-based encoders, facilitating energy-efficient 3D open-world understanding with a spike-driven nature.
4 Method
Our primary goal is to develop a spike-based encoder that accurately captures the geometric properties of 3D input data and efficiently achieves a unified representation for open-world 3D understanding with the spike-driven nature. To this end, we construct our triplet dataset $\{({D}^{t}_{1},{I}_{1}^{t},{T}_{1}^{t}),({D}^{t}_{2},{I}_{2},{T}_{2}^{t}),\cdots,({D}^{t}_{n},{I}_{n}^{t},{T}_{n}^{t})\}$, which consists of a 3D input ${D}^{t}_{i}$, an image ${I}_{i}$, and a text description ${T}_{i}$ at $t$ time step.
4.1 3D Input Representation
In this part, we present the 3D input representation, such as point clouds and event streams. Event streams, in particular, require special handling. We define them as $E_{i}=(x_{i},y_{i},t_{i},p_{i})$. Using a sliding window technique [23; 25], we convert event streams into an event cloud, formulated as:
$E_{i}=(x_{i},y_{i},z_{i})\quad\text{where}\quad z_{i}=\frac{t_{i}-t_{\text{min}}}{t_{\text{max}}-t_{\text{min}}},$
By doing so, we treat event streams as a distinct kind of spatio-temporal point cloud. This allows us to consider both point clouds and event streams as collections of points, denoted by ${D}^{t}=\{\mathcal{P},\mathcal{I}\}$. This includes voxel sets ${D}_{k}^{t}=\{\mathcal{P}_{k}^{t},\mathcal{F}_{k}^{t}\}$, where $\mathcal{P}_{k}^{t}\in\mathbb{R}^{3}$ represents the 3D coordinates and $\mathcal{I}_{k}^{t}\in\mathbb{R}^{D}$ indicate the features across $d$ channels at the time step $t$. Following this, we utilize our I-LIF spiking neuron to encode these 3D inputs into spatio-temporal spike trains, which are then transmitted to the spike encoder.
4.2 Multi-scale Triple Alignment
To develop a unified representation for open-world 3D understanding, we introduce a multi-scale triple alignment (MTA) framework that jointly optimizes correlation alignment across text, image, and 3D inputs. This framework integrates both semantic spike-text alignment and fine-grained spike-image alignment. Specifically, the overall architecture of SVL, illustrated in Fig. 1, comprises three encoders: (i) Text Encoder ($\mathcal{E}_{\theta}^{T}$): embeds text into text features $\mathcal{F}^{T}\in\mathbb{R}^{C}$; (ii) Spike-based Encoder ($\mathcal{E}_{\theta}^{S}$): transforms spike inputs into spike trains $\mathcal{F}^{S}\in\mathbb{R}^{T\times C}$. (iii) Image Encoder ($\mathcal{E}_{\theta}^{I}$): encodes images into image features $\mathcal{F}^{I}\in\mathbb{R}^{C}$. Here, $C$ represents the embedding dimension. These encoders collaboratively embed the triplet texts, spikes, and images into their respective feature spaces, facilitating comprehensive and fine-grained alignment across different modalities.
Semantic Spike-Text Alignment
To leverage the open-world recognition capabilities of the pre-trained CLIP model [13], we align the spike firing rate $\mathcal{F}^{S}/T$ with the text embeddings $\mathcal{F}^{T}$ obtained from CLIP, using a spike-text tuple $\mathcal{B}_{i}=\{{T}^{t}_{i},\mathcal{D}^{t}_{i}\}$ as input. The core idea is to bring the feature centroids of 3D instances and their corresponding text prompts closer together in the embedding space. To achieve this, we compute the InfoNCE loss [35] between the mean spike trains and the text features, as follows:
($\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{T})}=-\frac{1}{2|\mathcal{B}|}\sum_{i}^{|\mathcal{B}|}\log\frac{e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{j}}}+\log\frac{e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{x}_{j}\cdot y_{i}}},$ (5)
where $\mathbf{x}_{i}=\frac{\mathcal{F}^{S}/T}{||\mathcal{F}^{S}/T||2}$ and $\mathbf{y}_{i}=\frac{\mathcal{F}^{T}}{||\mathcal{F}^{T}||_{2}}$ represent the normalized spike and text features, respectively. The indices $i$ and $j$ are used for sampling, the dot product ($\cdot$) denotes cosine similarity between vectors, and $\tau$ is a learnable temperature parameter.
Fine-grained Spike-Image Alignment
A singular spike-text alignment fails to fully capture the semantic information embedded within both images and 3D data. To achieve a more comprehensive multimodal understanding, we further introduce an alignment between image and spike features. Specifically, we first employ the InfoNCE loss to align the image features, denoted as $\mathcal{F}^{I}$, with the average pulse signals, represented as $\mathcal{F}^{S}/T$. This alignment can be expressed as follows:
($\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{I})}=-\frac{1}{2|\mathcal{C}|}\sum_{i}^{|\mathcal{C}|}\log\frac{e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{i}}}{\sum_{j}^{|\mathcal{C}|}e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{j}}}+\log\frac{e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{a}_{j}\cdot b_{i}}},$ (6)
where $\mathcal{C}_{i}=\{{I}^{t}_{i},\mathcal{D}^{t}_{i}\}$ a spike-image tuple, $\mathbf{a}_{i}=\frac{\mathcal{F}^{S}/T}{||\mathcal{F}^{S}/T||2}$ and $\mathbf{b}_{i}=\frac{\mathcal{F}^{I}}{||\mathcal{F}^{I}||_{2}}$ represent the normalized spike and text features, respectively. However, this approach resulted in overly coarse alignment granularity, failing to account for the fine-grained and tightly coupled alignment between spikes and images. To address this, we incorporate the MSE loss on the basis of the InfoNCE loss to enhance the alignment granularity. The alignment objective between spike trains and images, which is formulated as follows:
$\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{MSE}}=\sum_{i}^{|\mathcal{C}|}||\mathcal{F}_{i}^{\textbf{S}}-\mathcal{F}_{i}^{\textbf{I}}||^{2},$ (7)
where $||\cdot||^{2}$ is the $\ell_{2}$ norm. Finally, we obtain the resultant total learning objective $\mathcal{L}_{\text{total}}$ as the combination of $\mathcal{L}_{(\textbf{S},\textbf{T})}$ and $\mathcal{L}_{(\textbf{S},\textbf{I})}$, where both alignments of semantic spike-text and fine-grained spike-image alignment are injected as:
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{T})}+\lambda_{2}\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{I})}+\lambda_{3}\mathcal{L}^{\text{MSE}}_{(\textbf{S},\textbf{I})},$ (8)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the influence of image features, text features, and spike trains. In this paper, the default setting of both is 1.
4.3 Re-parameterizable Vision-language Integration
After pretraining, the text prompt is applied to guide the learned spike-based representation to enable the following zero-shot transfer tasks. However, during inference, the traditional methods [14; 15] still require a high-parameter text encoder, which compromises the spike-driven nature and significantly impacts the deployment of neuromorphic hardware. To address this issue, we propose a novel structure called the re-parameterizable Vision-language module (Rep-VLI), which converts text embeddings into the weights of linear layers. It can be described as:
$W^{L}_{i}=e^{\tau}\mathcal{E}_{\theta}^{T}({T}_{i}^{t}),$ (9)
where $\mathcal{E}_{\theta}^{T}(\cdot)$ is the text encoder and $W_{i}$ is the re-parameterized layer weights on the $i$-th input text prompt ${T}_{i}^{t}$ at $t$ time step. Then we concatenate the re-parameterized weights $W_{i}$ by $K$ text prompts to $W^{L}\in\mathbb{R}^{K\times C}$. And the classification logits are calculated with the spike trains $\mathcal{F}^{S}$ and re-parameterized weights as:
$\text{logits}_{i}=\text{softmax}[\frac{W^{L}}{T}\mathcal{E}^{S}_{\theta}({D}^{t}_{i})],$ (10)
where $\mathcal{E}_{\theta}^{S}(\cdot)$ is the spike-based encoder, $T$ is the time step, and ${D}^{t}_{i}$ is the input 3D data. Rep-VLI overcomes the limitations of traditional high-parameter models by removing the need for a text encoder during inference. It re-parameterizes text into a lightweight format aligned with the spike-driven framework, ensuring efficient deployment on neuromorphic hardware while supporting diverse tasks like 3D open-world detection and segmentation [36; 37] with spike-based encoders.
5 Experiments
To validate that our spike-based 3D encoder learns robust visual representations via SVL, we evaluate it on diverse 3D open-world tasks, including zero-shot classification and visual question answering. The pretrained encoder is also fine-tunable for downstream tasks like 3D classification, segmentation, detection, and action recognition. This section details the experimental setup, including backbones, datasets, and implementation, followed by quantitative results and ablation studies on modality count, time steps, and loss functions. For further details about implementation and the datasets, please refer to Appendix E and Appendix F, respectively.
3D Backbone Networks
We evaluate three 3D backbones within our SVL framework: (i) Spike PointNet [26], a lightweight, energy-efficient spiking model for point clouds; (ii) E-3DSNN [5], a state-of-the-art model leveraging spike sparse convolution for sparse 3D inputs; and (iii) Spike PointFormer is our proposed Transformer-based SNN backbone for 3D data. Further details are provided in Appendix B. We use OpenCLIP’s [38] ViT-G/14 text encoder (frozen during pretraining) to define a shared feature space, into which the 3D modality is aligned. During inference, our Rep-VLI transforms text embeddings into compact weights, preserving the spike-driven nature of the above spike-based encoders.
Architecture Model Pre-train Method Input $T\times D$ Point+Text Param (M) Energy (mJ) Obj. M40. ANN PointCLIP [19] N/A Image N/A 25.5+57.3 24.9+20.3 1.9 20.2 PointNet [22] Openshape [16] Point N/A 3.47+202.5 20.1+71.7 24.4 74.9 Point-Bert [39] Point N/A 21.9+202.5 83.2+71.7 43.2 82.8 Sparseconv [40] Voxel N/A 5.3+202.5 0.61+71.7 31.7 78.8 Voxel N/A 41.3+202.5 2.13+71.7 43.4 83.4 Point-Bert [39] Ulip [14] Point N/A 21.9+227.8 81.2+80.6 34.9 69.6 Point-Bert [39] Ulip2 [15] Point N/A 21.9+202.5 83.7+71.7 50.6 84.7 SNN $\text{SpikeCLIP}^{*}$ [11] N/A Image 4$\times$1 9.5+22.8 10.6+0.41 0.5 5.1 Spike PointNet [25] SVL (Ours) Point 1$\times$4 3.57 0.27 24.9 76.3 Spike PointFormer Point 1$\times$4 22.1 9.4 43.4 83.1 E-3DSNN-T [5] Voxel 1$\times$4 2.1 0.04 33.6 79.6 E-3DSNN-S [5] Voxel 1$\times$4 3.5 0.09 36.4 81.3 E-3DSNN-L [5] Voxel 1$\times$4 17.7 0.64 43.9 84.6 E-3DSNN-H [5] Voxel 1$\times$4 46.7 0.79 47.0 85.4
Table 1: 3D Zero-shot classification results on the large-scale Objaverse-LVIS (Obj.) [41] and ModelNet40 (M40.) [24] datasets. "*" denotes self-implementation results with open-source code. "Energy" denotes the estimated energy consumption, following [5; 42]; further details are provided in Appendix D. "Point+Text" denotes the parameters of the point encoder and the text encoder.
5.1 3D Open-world Understanding
Zero-shot classification
We evaluate the zero-shot classification performance of our models on the widely-used ModelNet40 [24] and the larger, more challenging Objaverse-LVIS [41]. Compared to other benchmarks, Objaverse-LVIS offers broader class coverage and a long-tailed distribution, providing a more realistic evaluation of open-world 3D understanding [16]. As shown in Tab. 1, our SVL-based E-3DSNN achieves 85.4% accuracy on ModelNet40 with only 17.7M parameters, outperforming both ANN and SNN baselines. This demonstrates SVL’s effectiveness in enhancing both accuracy and efficiency.
Specifically, compared to OpenShape and ULIP, our model achieves 85.4% accuracy (vs. 83.4% and 69.6%), consumes only 0.79 mJ of energy (vs. 161.8 mJ and 73.8 mJ), and uses fewer parameters (17.7M vs. 41.3M and 21.9M). It also delivers a 15.8% accuracy gain over ULIP-based Point-BERT [14] while consuming just 11.4% of the energy. On Objaverse-LVIS, our model performs comparably to ULIP-2 [15] but with a 204$\times$ energy efficiency advantage. This is enabled by our Rep-VLI module, which reparameterizes text embeddings into compact, spike-compatible weights for zero-shot inference, preserving the spike-driven nature of the encoder. Compared to prior SNN approaches such as SpikeCLIP [11], SVL substantially improves the visual representation capacity of spike-based encoders in zero-shot 3D tasks.
Method Vision Encoder LLM Input S.-BERT SimCSE B-1. R-L. MET. InstructBLIP-13B [43] ViT [33] Vicua [44] Image 45.90 48.86 4.65 8.85 13.23 LLaVA-13B [45] Image 46.37 45.90 4.02 8.15 12.58 PointLLM-13B [12] PointBert [39] Vicua [44] Point 47.91 49.12 3.83 7.23 12.26 SVL-13B (Ours) Spike PointFormer Point 44.87 45.91 3.77 6.85 12.25 SVL-13B (Ours)* Spike PointFormer Point 47.80 47.08 11.45 14.69 16.40 Human N/A N/A N/A 100.00 100.00 100.00 100.00 100.00
Table 2: 3D object captioning results on Objaverse-LVIS. "*" indicates SVL-13B is prompted for shorter captions with no more than 20 words. The evaluation utilizes a range of metrics, including Sentence-BERT, SimCSE, BLEU-1, ROUGE-L, and METEOR.
Generative 3D Object Captioning and Open-world Question Answering
We combine the SVL-trained Spike PointFormer with a language model [44] via the LLaVA framework [45] for multimodal pre-training and fine-tuning (see Appendix B). On the 3D object captioning benchmark, prompted with “Describe this 3D model in detail,” our SVL-13B achieves performance comparable to state-of-the-art ANN methods (Tab. 2). Semantic metrics (Sentence-BERT, SimCSE) confirm strong alignment with human references. Notably, as the first SNN-based method in 3D captioning, SVL-13B achieves comparable annotation quality compared to PointLLM. We also evaluate on 3D question answering. As shown in Fig. 2, the model effectively interprets shape, material, function, and context, including visual and functional cues, while demonstrating commonsense reasoning. Despite lacking dense textures, SVL-13B achieves strong perception-language alignment, comparable to ANN models across diverse object types.
5.2 3D Downstream Tasks
3D Classification, Segmentation, and Detection
We first fine-tuned our models on 3D classification datasets such as ModelNet40 [24] and ScanObjectNN [46] to evaluate the 3D visual representation capabilities acquired through SVL pretraining. As shown in Tab. 3, the SVL pretraining significantly enhances performance, with the E-3DSNN [5] and the Spike PointNet [25] architecture achieving improvements of 1.9% and 1.0%, respectively, on ModelNet40. On the more challenging ScanObjectNN dataset, the Spike PointNet architecture demonstrates a substantial accuracy increase, rising from 70.0% to 76.1%. Subsequently, we extended our fine-tuning experiments to datasets such as Semantic KITTI [47] and KITTI [48]. As illustrated in Tab. 2, our SVL pretraining delivers marked improvements in both 3D segmentation and detection tasks, with the E-3DSNN [5] exhibiting performance gains of 1.1% and 1.2%, respectively.
Human Action Recognition
We further fine-tune our SVL-pretrained spike-based encoder on DVS datasets, including DVS128 Gesture [49] and DVS Action [50], to assess spatiotemporal feature extraction. During pretraining, the I-LIF time step was set to 1 for efficiency, then increased to 6 during evaluation to better capture temporal dynamics. We adopt the point-based method from [23] for efficient DVS data processing (see Section 4.1). As shown in Tab. 4, SVL-pretrained E-3DSNN and Spike Point improve by 2.1% and 1.6% on DVS Action and DVS128 Gesture, respectively, indicating strong scalability and temporal modeling ability of SVL-trained SNNs.
Architecture Model Input Param (M) Energy (mJ) $T\times D$ ModelNet40 ScanObjectNN ANN PointNet [22] Point 3.27 2.02 N/A 89.2 68.2 PointNet + ULIP [14] Point 3.47 2.34 N/A 92.1 72.1 Pointformer [51] Point 4.91 5.1 N/A 92.8 81.3 SNN P2SResLNet [26] Point 14.3 - $4\times 1$ 88.7 81.2 SpikingPointNet [52] Point 3.47 0.91 $16\times 1$ 88.6 66.6 Spike PointNet [25] Point 3.47 0.24 $1\times 4$ 88.2 70.0 Spike PointNet + SVL Point 3.47 0.27 $1\times 4$ 90.1 ($\uparrow 1.9$) 76.1 ($\uparrow 6.1$) E-3DSNN-S [5] Voxel 3.27 0.02 $1\times 4$ 91.7 78.7 E-3DSNN-S + SVL Voxel 3.27 0.02 $1\times 4$ 92.7 ($\uparrow 1.0$) 80.9 ($\uparrow 2.2$) E-3DSNN-L [5] Voxel 17.7 0.26 $1\times 4$ 91.2 80.2 E-3DSNN-L + SVL Voxel 17.7 0.31 $1\times 4$ 93.7 ($\uparrow 2.5$) 83.0 ($\uparrow 2.8$)
Table 3: 3D Downstream Tasks: 3D classification results on ModelNet40 (M-40) [24] and ScanObjectNN (Scan-O) [46].
Architecture Method $T\times D$ KITTI AP-E (%) Semantic KITTI mIoU (%) DVS Action Acc. (%) DVS128 Gesture Acc. (%) ANN E-3DANN [5] N/A 89.4 69.4 - - PointNet [22] N/A - 14.6 75.1 95.3 SNN E-3DSNN [5] $1\times 4$ 89.6 68.5 - - E-3DSNN + SVL $1\times 4$ 90.7 ($\uparrow 1.1$) 69.7 ($\uparrow 1.2$) - - Spike PointNet [25] $1\times 4$ / $6\times 4$ 12.1 78.4 96.9 Spike PointNet + SVL $1\times 4$ / $6\times 4$ 15.6 ($\uparrow 2.1$) 80.5 ($\uparrow 2.1$) 98.5 ($\uparrow 1.6$)
Table 4: 3D Downstream Tasks: 3D segmentation, detection, and human action recognition results on KITTI [48], Semantic KITTI [47], DVS Action [50], and DVS128 Gesture [49]. Moreover, for the DVS dataset, we adopt a pre-training timestep of $1\times 4$, consistent with other datasets, and a fine-tuning timestep of $6\times 4$.
5.3 Ablation Study
The Effectiveness of Our MTA
An ablation study was conducted to examine the impact of different loss function combinations during our multi-scale triple alignment (MTA). Specifically, we compared performance with and without the semantic spike-text alignment $(e.g.,\mathcal{L}_{(\mathbf{S},\mathbf{T})}^{\text{NCE}})$ and fine-grained spike-image alignment $(e.g.,\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{NCE}}$, $\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{MSE}})$. As shown in Tab. 6, the ablation study highlights the importance of combining loss functions for optimal performance. In the absence of any loss functions, the model only gets 0.5% accuracy on the large-scale Objaverse-LVIS and 5.1% on ModelNet40. Introducing spike-image alignment yields a significant improvement, while the inclusion of semantic spike-text alignment alone demonstrates limited effectiveness. The highest performance is attained when all three loss functions, including the MSE-based fine-grained alignment, are employed, achieving 33.6% accuracy on the large Objaverse-LVIS [41] and 79.6% on ModelNet40 [24]. These findings underscore the synergistic relationship between semantic and fine-grained alignment in enhancing the model’s representational capabilities, showing the effectiveness of our MTA module.
Different Time Steps and Firing Bits Analysis
We systematically evaluated the impact of varying $T$ and $D$ on both SVL pretraining and downstream task fine-tuning, as shown in Tab. 6. For pretraining, we observed that increasing the number of $T$ does not enhance performance, while it adversely affects inference time and significantly increases power consumption. Specifically, after our SVL pretraining, we tested the effect of different $T$ and $D$ on the zero-shot classification accuracy of large-scale Objaverse-LVIS [41]. For instance, with $D=1$, extending $T$ from 2 to 4 resulted in a 0.2% increase in accuracy, accompanied by a doubling of power consumption. Conversely, we found that expanding $D$ while maintaining a fixed $T$ not only improves performance but also reduces power consumption. For fine-tuning, increasing the number of time steps enhances performance but comes at the cost of longer inference time and higher power consumption. In summary, increasing $T$ enhances fine-tuning performance but requires balancing to limit power consumption and inference time, while expanding $D$ offers a solution for improving performance and reducing power demands.
Table 5: Ablation study of the pretrain timesteps.
Method $T\times D$ Power (mJ) Obj. (%) M40. (%) $\text{ANN}^{*}$ N/A 0.13 34.1 81.3 SNN $1\times 2$ 0.02 32.9 78.5 $2\times 1$ 0.03 32.7 78.0 $2\times 2$ 0.08 33.9 80.5 $1\times 4$ 0.04 33.6 79.6 $4\times 1$ 0.10 32.9 78.6
Table 6: Ablation study of MTA.
$\mathcal{L}_{(\mathbf{S},\mathbf{T})}^{\text{NCE}}$ $\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{NCE}}$ $\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{MSE}}$ Obj. M40. 0.5 5.1 24.8 73.1 21.9 70.1 31.7 77.8 33.6 79.6
Figure 2: Dialogues between SVL-13B and a human user. The dialogues show SVL’s ability to understand point clouds’ shapes, appearances, functionalities, etc. Additionally, SVL demonstrates abilities to respond to human instructions with common sense, avoiding biases.
6 Conclusion
In this work, we introduce SVL, a spike-based vision-language pretraining framework that equips SNNs with open-world 3D understanding while preserving their inherent energy efficiency. By integrating Multi-scale Triple Alignment (MTA) and a Reparameterizable Vision-Language Integration (Rep-VLI) module, SVL bridges the gap between the low-power advantages of SNNs and the strong generalization capabilities of vision-language models. Comprehensive evaluations across zero-shot 3D classification, semantic segmentation, and human action recognition demonstrate that SVL consistently outperforms prior SNN-based approaches and even rivals state-of-the-art ANNs, all with significantly lower computational cost. Notably, SVL enables SNNs to perform open-world 3D question answering, marking a milestone in multimodal representation learning for spike-based systems. In summary, SVL enables SNNs to acquire powerful, transferable multimodal representations, paving the way for advanced 3D understanding in resource-constrained environments and laying the groundwork for large-scale multimodal learning in neuromorphic computing.
Appendix
Appendix A Backpropagation process of I-LIF
There exist two primary methods of training high-performance SNNs. One way is to discretize ANN into spike form through neuron equivalence [53], i.e., ANN-to-SNN conversion, but this requires a long simulation time step and boosts the energy consumption. We employ the direct training method [29, 54] and apply surrogate gradient training.
Then in this section, we introduce the training process of SNN gradient descent and the parameter update method of spatio-temporal backpropagation (STBP) [29, 55, 56]. SNNs’ parameters can be taught using gradient descent techniques, just like ANNs, after determining the derivative of the generation process. Moreover, the accumulated gradients of loss $\mathcal{L}$ with respect to weights $\mathbf{w}$ at layer $\ell$ can be calculated as:
$\small\frac{\partial\mathcal{L}}{\partial{W}^{\ell}}=\sum_{t=1}^{T}\frac{\partial\mathcal{L}}{\partial{s}^{\ell+1}[t]}\frac{\partial{s}^{\ell+1}[t]}{\partial{u}^{\ell+1}[t]}\left(\frac{\partial{u}^{\ell+1}[t]}{\partial{w}^{\ell}}+\sum_{\tau<t}\prod_{i=t-1}^{\tau}\left({\frac{\partial{u}^{\ell+1}[i+1]}{\partial{u}^{\ell+1}[i]}}+{\frac{\partial{u}^{\ell+1}[i+1]}{\partial{s}^{\ell+1}[i]}}{\frac{\partial{s}^{\ell+1}[i]}{\partial{u}^{\ell+1}[i]}}\right)\frac{\partial{u}^{\ell+1}[\tau]}{\partial{W}^{\ell}}\right),$ (11)
where ${s}^{\ell}[t]$ and ${u}^{\ell}[t]$ represent the binary and membrane potential of the neuron in layer $\ell$, at time $t$. Moreover, notice that $\frac{\partial{s}^{\ell}[t]}{\partial{u}^{\ell}[t]}$ is non-differentiable. To overcome this problem, [29] propose the surrogate function to make only the neurons whose membrane potentials close to the firing threshold receive nonzero gradients during backpropagation. In this paper, we use the rectangle function, which has been shown to be effective in gradient descent and may be calculated by:
$\frac{\partial{s}^{\ell}[t]}{\partial{u}^{\ell}[t]}=\frac{1}{a}\operatorname{sign}\left(\left|{u}^{\ell}[t]-\vartheta\right|<\frac{a}{2}\right),$ (12)
where $a$ is a defined coefficient for controlling the width of the gradient window, and is set to 1 in our paper.
Appendix B Architecture Details
In this section, we present the detailed architectural designs of E-3DSNN, Spike PointNet and Spike PointFormer, outlining their core components, network configurations, and the specific adaptations made to enable efficient analysis of 3D point cloud within the spiking neural network framework.
E-3DSNN [5]
are realized by adjusting the number of blocks and channels across stages to balance model size and performance. As shown in Tab. 7, the architecture scales from lightweight (E-3DSNN-T) to high-capacity (E-3DSNN-H) models, with corresponding changes in parameters and feature dimensions.
Types Blocks Channels Param. (M)
E-3DSNN-T [1, 1, 1, 1] [16, 32, 64, 128] 1.8
E-3DSNN-S [1, 1, 1, 1] [24, 48, 96, 160] 3.2
E-3DSNN-L [2, 2, 2, 2] [64, 128, 128, 256] 17.3
E-3DSNN-H [2, 2, 2, 2] [96,192,288,384] 46.5
Table 7: Architecture details of E-3DSNN [5]
Spike PointNet [25]
is the first spiking neural network specifically designed for efficient deep learning on 3D point clouds. It leverages the sparse and event-driven nature to achieve high accuracy with few parameters and low power consumption. This makes it particularly well-suited for deployment in energy-constrained or real-time 3D perception scenarios.
Spike PointFormer
is our proposed Transformer-based SNN backbone for point cloud encoding. Built upon PointBERT [39], we replace ReLU with I-LIF to introduce spiking dynamics. Specifically, the point cloud is first downsampled and grouped into local regions. These are passed through the SNN-adapted encoder to extract local features. Then, Spike PointFormer stacks $L$ spike-driven Transformer layers, enabling global feature interaction. Together, Spike PointFormer can be formulated as:
$x={\rm{KNN}}(\rm{FPS}(P)),$ $P\in\mathbb{R}^{B\times N\times 3},\quad x\in\mathbb{R}^{B\times N^{\prime}\times K\times 3}$ (13)
$f_{0}={\rm{EMP}}({\rm{MLP}}({\mathcal{SN}}\left(x\right))),$ $f_{0}\in\mathbb{R}^{B\times N^{\prime}\times D}$ (14)
$f_{l}={\rm{SDT}}(f_{l-1})+f_{l-1},$ $f_{l}\in\mathbb{R}^{B\times N^{\prime}\times D},l=1...L$ (15)
where $P\in\mathbb{R}^{B\times N\times 3}$ is the input point cloud with batch size $B$, point number $N$. ${\rm FPS}(\cdot)$ and ${\rm KNN}(\cdot)$ denote Farthest Point Sampling and K-nearest Neighbor Grouping. ${\mathcal{SN}}(\cdot)$ represents spike neuron; ${\rm MLP}(\cdot)$ consists of multiple convs; ${\rm EMP}(\cdot)$ denotes Element-wise Max Pooling. ${\rm SDT}(\cdot)$ denotes Spike-Driven Transformer layer. $f_{0}$ and $f_{l}$ are intermediate features with dimension $D$.
Notably, to better align with SNN principles, we remove the softmax in the spike-driven transformer layer and re-arrange the computation order of Query, Key and Value to achieve sparse and energy-efficient attention. In particular, the spike-driven attention mechanism can be formulated as:
$\operatorname{SDA}\left(Q,K,V\right)=\mathcal{SN}\left(\mathcal{SN}\left(Q\right)\odot\mathcal{SN}\left(K\right)^{\mathrm{T}}\right)\odot\mathcal{SN}\left(V\right),$ (16)
where $\operatorname{SDA}(\cdot)$ denotes the spike-driven attention, and $\odot$ represents the dot-product operation. $\mathcal{SN}(\cdot)$ represents spike neurons referred to I-LIF.
Appendix C Open-world Multimodel Learning Details
Figure 3: Architecture details of open-world multimodel learning.
In this section, we present the details of how to perform open-world multimodal learning after pretraining with the SVL model. Our primary goal is to effectively leverage the capabilities of the pretrained LLM and the spike-based encoder pretrained with SVL. The network architecture is shown in Fig. 3. We select vicuna [44] as our LLM $\mathcal{F}_{\theta}(\cdot)$, parameterized by $\theta$, use the Spike Pointformer $\mathcal{E}^{S}_{\theta}(\cdot)$ pretrained with SVL as the spiking visual encoder, and the projector $\mathcal{P}_{\theta}(\cdot)$.
For the input 3D data $D^{t}_{k}$, we first utilize the pretrained spike-based encoder $\mathcal{E}^{S}_{\theta}(\cdot)$ to provide 3D spatiotemporal visual features. Subsequently, a simple linear layer $\mathcal{P}_{\theta}(\cdot)$ is employed to connect the 3D spatiotemporal visual features to the word embedding space. Specifically:
$V^{t}=\mathcal{P}_{\theta}(\mathcal{E}^{S}_{\theta}(D^{t}_{k})),$ (17)
where $V^{t}$ is the vison tokens at the $t$ time step. Then we concatenate it with the text tokens $I^{t}$ obtained after tokenization and send the combined features to the LLM $\mathcal{F}_{\theta}(\cdot)$.
$R^{t}=\mathcal{F}_{\theta}(I^{t};V^{t}),$ (18)
where $R^{t}$ is the output response or logits. Our training is divided into two stages. In the first stage, we train the projector while freezing the LLM and the spike-based encoder. In the second stage, we train the LLM and the projector.
Table 8: Qualitative comparisons. We show the qualitative results of models on the ScanNet [57]. Our SVL-13B can understand 3D semantics and respond to prompts effectively comparable to other ANN-based models.
SVL-13B (Ours) (The outputs for ScanNet-Scene0024_02 are shown below.)
[Uncaptioned image] This 3D model depicts a traditional-style house with a tasteful aesthetic. The house features a rich brown color on its exterior walls, giving it a warm and welcoming appearance. It has a distinctive architectural design, with a slanted roof that is typical of traditional homes. The interior reflects a beautiful contrast with lighter-colored walls, providing a homely and comfortable ambiance. This model can be used in architectural designing, virtual reality games, or various design projects.
Table 9: Qualitative comparisons. We show the qualitative results of models on ModelNet40 [24] and Objaverse [41]. Our SVL-13B can understand 3D semantics and respond to prompts effectively comparable to other ANN-based models.
       Samples         [Uncaptioned image][Uncaptioned image]         [Uncaptioned image][Uncaptioned image]
       Ground Truth         Laptop         A cartoon black monster like a dragon
       Prompt         What is this?         Briefly caption this 3D model.
       InstructBLIP [43]         symbol letter l         a black lizard with a sharp tooth in a dark room
       LLaVA [45]         A small, grainy, black and white letter j.         A 3D model of a dark, menacing dragon.
       3D-LLM [58]         -         A black and white tiger with long legs, standing on its hind leg.
       Point-Bind LLM [59]         This is a laptop computer.         The 3D model features a large, ornate gargoyle with a horned helmet, sitting on top of a building.
       PointLLM [12]         The 3D model represents a notebook computer, typically a laptop.         The 3D model depicts a menacing black dragon, with its mouth opened wide revealing a row of sharp teeth.
       SVL-13B (Ours)         This is a 3D model of a laptop.         This is a 3D model of a toy dinosaur, which stands upright on its hind legs. It has a spiked back, reflecting its distinctive defense mechanism.
Appendix D Theoretical Energy Consumption
In our SVL framework, Rep-VLI can transform the text embeddings into tiny weights during inference. Additionally, the framework can convert matrix multiplication into sparse addition, which can be implemented as addressable additions on neuromorphic chips. In the first coding layer, convolution operations act as Multiply-Accumulate (MAC) operations that convert analog inputs into spikes, similar to direct coding-based SNNs [60]. Similarly, in the final layer, logit calculations also perform MAC operations. In contrast, in the SNN architecture, the convolution (Conv) or fully connected (FC) layers transmit spikes and execute Accumulation (AC) operations to accumulate weights for postsynaptic neurons. Hence, the inference energy cost for our SVL framework can be expressed as follows:
$E_{total}=E_{MAC}\cdot(FL_{conv}^{1}+FL_{conv}^{VLI})+E_{AC}\cdot T\sum_{n=2}^{N}FL_{conv}^{n}\cdot fr^{n},$ (19)
where $N$ and $M$ represent the total number of sparse spike convolutions, and $E_{MAC}$ and $E_{AC}$ are the energy costs associated with MAC and AC operations, respectively. The variables $fr^{m}$, $fr^{n}$, $FL_{conv}^{n}$, and $FL_{fc}^{m}$ denote the firing rate and FLOPs of the $n$-th sparse spike convolution layer. Previous SNN studies [61, 62, 54, 63] assume a 32-bit floating-point implementation in 45nm technology, with $E_{MAC}$ = 4.6 pJ and $E_{AC}$ = 0.9 pJ for various operations.
Additionally, batch normalization (BN) operations can be fused into the convolutional layers, further reducing computation overhead. Since Rep-VLI eliminates the text encoder during inference, layer normalization (LN) layers are also unnecessary, simplifying the architecture and lowering energy consumption. These design choices ensure that our framework is both energy-efficient and optimized for neuromorphic deployment.
Appendix E Implementation Details
The hyperparameters for SVL pretraining are presented in Tab. 10. The hyperparameters for SVL fine-tuning on 3D point clouds are detailed in Tab. 11, and those for SVL fine-tuning on DVS are outlined in Tab. 12.
Table 10: Hyper-parameters for SVL pretraining.
Hyper-parameter E-3DSNN-T/S/L/H Spike PointNet Spike PointFormer
Timestep (Training/Inference) $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$
Epochs 250 250 250
Batch size 4096 1024 1024
Optimizer AdamW AdamW AdamW
Base Learning rate $32e-$ 3 $32e-$ 3 $33e-$ 3
Learning rate decay Cosine Cosine Cosine
Warmup eopchs 10 10 10
Weight decay $41e-$ 4 $41e-$ 4 0.1
Table 11: Hyper-parameters for SVL Finetuning on 3D point cloud.
Hyper-parameter ModelNet40 ScanObjectNN KITTI SemanticKITTI
Timestep (Training/Inference) $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$
Epochs 300 250 100 80
Batch size 64 64 96 64
Optimizer AdamW AdamW AdamW AdamW
Base Learning rate $45e-$ 4 $35e-$ 3 $21e-$ 2 $32e-$ 3
Learning rate decay Onecycle Onecycle Onecycle Onecycle
Table 12: Hyper-parameters for SVL Finetuning on DVS datasets.
Hyper-parameter DVS Action DVS128 Gesture
Timestep (Training/Inference) $1\times 4$/ $6\times 4$ $1\times 4$/ $6\times 4$
Epochs 250 250
Batch size 4096 1024
Optimizer AdamW AdamW
Base Learning rate $32e-$ 3 $32e-$ 3
Learning rate decay Cosine Cosine
Warmup eopchs 10 10
Weight decay $41e-$ 4 $41e-$ 4
Appendix F Datasets
[…]
The Objaverse dataset, which includes Objaverse-LVIS [41] as a subset, is currently the largest 3D dataset. Objaverse-LVIS is a significant part of the Objaverse dataset, containing 46,832 annotated shapes across 1,156 LVIS categories. This extensive collection of 3D shapes provides a rich resource for researchers and practitioners in the field of computer vision and 3D modeling.
The large KITTI dataset [64] contains 7481 training samples, 3717 of which constitute trainsets and 3769 of which constitute validation sets. E-3DSNN is evaluated as backbones equipped with VoxelRCNN Head In detection [65]. To execute our model, we uses OpenPCDet [66] that is transformed into a spiking version by us. After being divided into regular voxels, raw point clouds are input to our 3DSNN on KITTI [48]. The point clouds are clipped to ranges of $[-0.7\text{m},0.4\text{m}]$ for the X-axis, $[-40\text{m},40\text{m}]$ for the Y-axis, and $[-3\text{m},1\text{m}]$ for the Z-axis followed by voxelization with a resolution of $(0.05\text{m},0.05\text{m},0.1\text{m})$. The Average Precision (AP) calculated by 11 recall positions for the Car class is used as the evaluation metrics.
[…]
The DVS Action dataset [50] comprises 10 actions performed by 15 subjects within 5s, which is recorded by DVS camera in an empty office. DVS is a vision sensor [50] that can records a sequence of tuples $[t,x,y,p]$ for each event streams. Among them, $t$ represents the timestamp of the event, $(x,y)$ represents the event’s pixel coordinates and $p$ represents the polartity of the event.
The DVS128 Gesture dataset [49] contains 1,342 instances across 11 different hand and arm gestures, which are performed by 29 subjects under 3 distinct lighting conditions in 122 trials. They are captured by DVS128 camera, a DVS with $128\times 128$ pixel resolution.
Appendix G Limitations
One of the limitations of our work is exploring scalability of E-3DSNN. The largest model reported is E-3DSNN-H with 46.7M parameters. Future work will explore the performance and computational efficiency trade-offs of SVL at larger parameter scales.
The experimental results presented in this paper are reproducible. Detailed explanations of model training and configuration are provided in the main text and supplemented in the appendix. Our codes and models will be made available on GitHub after review.